if one member is chosen at random from a club that has 12 girls and 9 boys, what is the probability that a girl will be selected?

Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

12/21=4/7=.5714 or 57.14 % chance of selecting a girl.

Answer by stanbon(75887)

if one member is chosen at random from a club that has 12 girls and 9 boys, what is the probability that a girl will be selected?
-------------------------------
P(select a girl) = 12/21 = 4/7
========================================
Cheers,
Stan H.